how do you solve (x+2)^2/9- (y+8)^2/4= 1
Answer by lwsshak3(11628)

how do you solve (x+2)^2/9- (y+8)^2/4= 1
This is an equation for a hyperbola with horizontal transverse axis of the standard form:
(x-h)^/a^2-(y-k)^2/b^2=1, with (h,k) being the (x,y) coordinates of the center.
..
For given equation:
center: (-2,-8)
a^2=9
a=√9=3
b^2=4
b=√4=2
c^2=a^2+b^2=9+4=13
c=√13≈3.6
Vertices: (-2±a,-8)=(-2±3,-8)=(-5,-8) and (1,-8)
Foci: (-2±c,-8)=(-2±√13,-8)=(-2±3.6,-8)=(-5.6,-8) and (1.6,-8)